# Towards Automatic Evaluation and High-Quality Pseudo-Parallel Dataset Construction for Audio Editing: A Human-in-the-Loop Method

### v1

@@ Appendix A Appendix
$\bullet$ 5 - About the Same/Even Better: The edited audio is on par with or even better than the original in terms of intelligibility, fluency, and background noise.
$\bullet$ 4 - Slightly Worse: At most one aspect (e.g., mild mechanical artifact or minor increase in noise) is slightly degraded, but overall listening quality remains unaffected.
$\bullet$ 3 - Noticeably Worse: At least two aspects show moderate degradation, or one aspect is clearly degraded. Artifacts such as robotic tone, discontinuities, or increased noise may affect understanding in some segments but remain tolerable.
$\bullet$ 2 - Much Worse: Significant degradation in at least two aspects, with noticeable artifacts severely affecting naturalness and intelligibility.
$\bullet$ 1 - Badly Damaged: Severe issues across all aspects, such as overwhelming noise, harsh distortions, or unintelligible content, rendering the audio difficult to comprehend.
@@
$\bullet$ 5 - Excellent Match: The edit perfectly matches the described intent. For Addition, the added event feels natural and contextually appropriate. For Deletion, the targeted event is successfully removed without leaving any noticeable residue. For Replacement, the substituted content is fluent and coherent, maintaining the overall auditory experience.
$\bullet$ 4 - Good Match: Overall alignment is correct, but there may be minor imperfections, such as slightly abrupt additions or faint remnants after deletion.
$\bullet$ 3 - Partial Match: The general category or type of edit is correct, but details may be awkward or somewhat misaligned with the description.
$\bullet$ 2 - Poor Match: The edit loosely corresponds to the intent, but contains semantic inaccuracies or unnatural execution that hampers comprehension.
$\bullet$ 1 - Mismatch: The edited content fails to reflect the described intent, with missing, incorrect, or misleading modifications.
@@
$\bullet$ 5 - Perfectly Faithful: Unedited segments are completely preserved. Prosody, order, and audio style remain fully consistent with the original.
$\bullet$ 4 - Mostly Faithful: Slight variation in loudness or subtle stylistic shifts may occur, but prosody and content are largely preserved.
$\bullet$ 3 - Partially Faithful: Noticeable changes in rhythm or sequence are present, though the overall style remains similar. Minor alterations in the unedited content may be perceptible.
$\bullet$ 2 - Low Faithfulness: Major changes in prosody, ordering, or style, with only a few elements traceable to the original.
$\bullet$ 1 - Not Faithful: The edited audio bears little or no resemblance to the original; unedited content is altered significantly or missing altogether.
Appendix C B. Details of Pseudo-Parallel Audio Editing Dataset Construction.
Inspired by AUDIT and InstructME, we also adopt an audio mixing strategy to construct our Pseudo-Parallel Audio Editing Dataset. The goal of the dataset construction process is to generate a set of audio pairs comprising (A, B, C, A + B, A + C), which serves as the foundation for subsequent audio editing tasks. For audio types A, we collected over 45,000 audio prompts from the AudioCaps dataset (Kim et al. 2019). For audio types B and C, we select prompts corresponding to common single-audio events, covering eight broad categories—NaturalSounds, Transportation, IndoorActivities, OutdoorActivities, Animals, Music, Speech and HumanSounds—spanning over 100 distinct sound types. By pairing these prompts in all possible two-way combinations, we obtained 4950 pairs of audio prompts. Then, for each audio prompt pair (B, C), we randomly select 20 audio prompts from the 45,000+ A audio prompts to form combinations, denoted as (A, B, C).
We utilize TangoFlux(Hung et al. 2024) to generate audio based on the (A, B, C) prompts, using a step of 50 and a guide scale of 4.5. The generation rules are as follows:
@@
Appendix D C. Filtering Strategy of High-Quality Pseudo-parallel Audio Editing Dataset
Table 5: Number of Samples After Each Round of Selection
@@
Here, we present the filtering strategy of constructing High-Quality Pseudo-Parallel Audio Editing Dataset. In AppendixC, the construction of the Pseudo-Parallel Audio Editing Dataset is presented, comprising 60,000 training instances for each of the three editing tasks: Addition, Deletion, and Replacement. We employ our proposed model, AuditEval, to evaluate the dataset across three dimensions(Quality, Relevance, and Faithfulness). Subsequently, based on the scores of these three dimensions, we perform three rounds of filtering on the dataset. The filtering criteria are detailed as follows:
  • Stage 1: Faithfulness-based filtering. All samples are sorted by their predicted faithfulness scores in descending order. We retain the top 30% of samples as determined by this ranking.
  • Stage 2: Quality thresholding. Among the retained samples, we further filter out those with predicted quality scores below 3.4.
@@
In practice, we use a quality threshold of 3.40, a faithfulness threshold of 3.2156/3.2813/3.2662 and a relevance threshold of 3.1592/3.4844/3.3441. Table5 illustrates the selection process, where the chosen thresholds approximately eliminate half of the data at each round.
Appendix E D. User Interface for Subjective Evaluation
To ensure consistency and ease of use during the manual scoring process, we designed a custom web-based interface for subjective evaluation. As shown in Figure 5, the interface presents the original audio, the edited audio, the original prompt, the edited prompt, and, if available, the edit instruction. Annotators are instructed to score each sample along three dimensions—Quality, Relevance, and Faithfulness—following the same evaluation criteria defined in the AuditScore dataset. This standardized interface streamlines the annotation workflow and helps reduce variability in human ratings.
@@
Appendix F E. Computational Experiments Details
@@
All experiments were conducted on a single consumer-grade GPU: NVIDIA GeForce RTX 4090. To ensure reproducibility, we fixed all sources of randomness with a global random seed set to 1984.
@@
The AuditEval model was trained using the Adam optimizer with a learning rate of $5\times 10^{-5}$. A learning rate scheduler ReduceLROnPlateau was employed with the following parameters: mode=’min’, factor=0.8, patience=5, and verbose=True.
The training was performed with a batch size of 256 for training and 64 for validation.
The model consists of three independent regression heads corresponding to the Quality, Relevance, and Faithfulness dimensions, trained for different numbers of epochs: Quality: trained for 50 epochs; Relevance: trained for 30 epochs; Faithfulness: trained for 35 epochs.
Appendix G F. Licenses for assets
@@
This work uses publicly available resources. AudioCaps and the audio editing inversion code are released under the MIT License. CLAP pretrained weights are under the CC0-1.0 License. All assets developed in this work—including datasets, checkpoints, training protocols, and code—will be released under the CC BY 4.0 License.
Appendix H G. Ethics Statement
@@
We employed professional annotators via a third-party commercial provider to ensure consistent and reliable subjective evaluations. Annotators were fairly compensated above local minimum wage in accordance with regional labor laws. They received clear instructions on task objectives, scoring criteria (Quality, Relevance, Faithfulness), and example cases prior to annotation. No personal or sensitive information was collected, and no individual-level data beyond task performance was recorded. The provider ensured fair labor practices, contractual compliance, and data confidentiality. As annotators were not research subjects but task workers evaluating system outputs, IRB approval was not required.
@@
Appendix I H. Evaluation scores Across Editing Operations (Add / Delete / Replace)

### v2

@@ Appendix A Details of Scoring Criteria Specification
  • 5- About the Same/Even Better: The edited audio is on par with or even better than the original in terms of intelligibility, fluency, and background noise.
  • 4- Slightly Worse: At most one aspect (e.g., mild mechanical artifact or minor increase in noise) is slightly degraded, but listening quality remains unaffected.
  • 3- Noticeably Worse: At least two aspects show moderate degradation, or one aspect is clearly degraded. Artifacts such as robotic tone, discontinuities, or increased noise may affect understanding in some segments but remain tolerable.
  • 2- Much Worse: Significant degradation in at least two aspects, with noticeable artifacts severely affecting naturalness and intelligibility.
  • 1- Badly Damaged: Severe issues across all aspects, such as overwhelming noise, harsh distortions, or unintelligible content, rendering the audio difficult to comprehend.
@@
  • 5- Excellent Match: The edit perfectly matches the described intent. For Addition, the added event feels natural and contextually appropriate. For Deletion, the targeted event is successfully removed without leaving any noticeable residue. For Replacement, the substituted content is fluent and coherent, maintaining the overall auditory experience.
  • 4- Good Match: Overall alignment is correct, but there may be minor imperfections, such as slightly abrupt additions or faint remnants after deletion.
  • 3- Partial Match: The general category or type of edit is correct, but details may be awkward or somewhat misaligned with the description.
  • 2- Poor Match: The edit loosely corresponds to the intent, but contains semantic inaccuracies or unnatural execution that hampers comprehension.
  • 1- Mismatch: The edited content fails to reflect the described intent, with missing, incorrect, or misleading modifications.
@@
  • 5- Perfectly Faithful: Unedited segments are completely preserved. Prosody, order, and audio style remain fully consistent with the original.
  • 4- Mostly Faithful: Slight variation in loudness or subtle stylistic shifts may occur, but prosody and content are largely preserved.
  • 3- Partially Faithful: Noticeable changes in rhythm or sequence are present, though the overall style remains similar. Minor alterations in the unedited content may be perceptible.
  • 2- Low Faithfulness: Major changes in prosody, ordering, or style, with only a few elements traceable to the original.
  • 1- Not Faithful: The edited audio bears little or no resemblance to the original; unedited content is altered significantly or missing altogether.
Appendix B Details of Pseudo-Parallel Audio Editing Dataset Construction.
Inspired by AUDIT and InstructME, we also adopt an audio mixing strategy to construct our Pseudo-Parallel Audio Editing Dataset. The goal of the dataset construction process is to generate a set of audio pairs comprising (A, B, C, A + B, A + C), which serves as the foundation for subsequent audio editing tasks. For audio types A, we collected over 45,000 audio prompts from the AudioCaps dataset Kim et al. [2019]. For audio types B and C, we select prompts corresponding to common single-audio events, covering eight broad categories—NaturalSounds, Transportation, IndoorActivities, OutdoorActivities, Animals, Music, Speech and HumanSounds—spanning over 100 distinct sound types. By pairing these prompts in all possible two-way combinations, we obtained 4950 pairs of audio prompts. Then, for each audio prompt pair (B, C), we randomly select 20 audio prompts from the 45,000+ A audio prompts to form combinations, denoted as (A, B, C).
We utilize TangoFlux Hung et al. [2024] to generate audio based on the (A, B, C) prompts, using a step of 50 and a guide scale of 4.5. The generation rules are as follows:
@@
Figure 3: Score prediction distributions on the large-scale Pseudo-Parallel Audio Editing Dataset, with addition, deletion, and modification operations shown from top to bottom.
Figure 4: Score prediction distributions on the large-scale Pseudo-Parallel Audio Editing Dataset, with addition, deletion, and modification operations shown from top to bottom.
Appendix C Filtering Strategy of High-Quality Pseudo-parallel Audio Editing Dataset
Table 5: Number of Samples After Each Round of Selection(ssl)
@@ Appendix C Filtering Strategy of High-Quality Pseudo-parallel Audio Editing Dataset
Table 6: Number of Samples After Each Round of Selection(llm)
Selection Stage Add Delete Replace
Original number 198000 198000 198000
After Round 1 99000 99000 99000
After Round 2 68551 76358 68522
After Round 3 30000 30000 30000
Here, we present the filtering strategy of constructing High-Quality Pseudo-Parallel Audio Editing Dataset. In AppendixC, the construction of the Pseudo-Parallel Audio Editing Dataset is presented, comprising 30,000 training instances for each of the three editing tasks: Addition, Deletion, and Replacement. We employ our proposed model, AuditEval, to evaluate the dataset across three dimensions, with the score distributions illustrated in Fig. 3 and Fig. 4. Subsequently, based on the scores of these three dimensions, we perform three rounds of filtering on the dataset. The filtering criteria are as follows:
  • Stage 1: Faithfulness-based filtering. All samples are sorted by their predicted faithfulness scores in descending order.
  • Stage 2: Quality thresholding. Among the retained samples, we further filter out those with predicted quality scores boardline.
@@
For AuditEval-ssl, we adopt a quality threshold of 3.40, faithfulness thresholds of 3.216 / 3.281 / 3.266, and relevance thresholds of 3.159 / 3.484 / 3.344 for the three successive filtering rounds. For AuditEval-llm, the corresponding thresholds are set to 3.70 for quality, 3.578 / 3.641 / 3.453 for faithfulness, and 3.500 / 3.609 / 3.469 for relevance.
Tables 5 and 6 illustrate the multi-round selection process for the two evaluators, respectively. In each round, the selected thresholds are designed to remove approximately half of the remaining samples, progressively retaining higher-quality pseudo-parallel data.
Appendix D User Interface for Subjective Evaluation
@@
To ensure consistency and ease of use during the manual scoring process, we designed a custom web-based interface for subjective evaluation. As shown in Figure 5, the interface presents the original audio, the edited audio, the original prompt, the edited prompt, and, if available, the edit instruction. Annotators are instructed to score each sample along three dimensions: Quality, Relevance, and Faithfulness, following the same evaluation criteria defined in the AuditScore dataset. This standardized interface streamlines the annotation workflow and helps reduce variability in human ratings.
Appendix E Computational Experiments Details
@@
All experiments were conducted on a single NVIDIA H800 GPU. To ensure reproducibility, we fixed all sources of randomness with a global random seed set to 1984.
@@
The AuditEval-ssl model was trained using the Adam optimizer with a learning rate of $5\times 10^{-5}$. We employed a ReduceLROnPlateau scheduler with mode=’min’, factor=0.8, and patience=5. Training was conducted with a batch size of 256 for training and 64 for validation. The model comprises three independent regression heads for Quality, Relevance, and Faithfulness, which were trained for 50, 30, and 35 epochs, respectively.
AuditEval-llm was trained within the ms-swift framework as a regression-based sequence classification model that predicts three continuous scores corresponding to Quality, Relevance, and Faithfulness (num_labels=3, problem_type=regression), with chat templates disabled to ensure a unified input format. We adopted LoRA-based parameter-efficient fine-tuning, applying LoRA to all linear layers with a rank of 8 and a scaling factor (alpha) of 32, while freezing the ViT backbone. The model was trained for 6 epochs using the Adam optimizer with a learning rate of $1\times 10^{-4}$ and a warmup ratio of 0.05, under mixed-precision training with bfloat16. The per-device batch size was set to 8 for training and 2 for evaluation, with gradient accumulation over 8 steps and a maximum input sequence length of 4096 tokens.
Appendix F Licenses for assets
@@
All assets developed in this work, including datasets, checkpoints, training protocols, and code, will be released under the CC BY 4.0 License.
Appendix G Ethics Statement
@@
We employed professional annotators via a third-party commercial provider to ensure consistent and reliable subjective evaluations. Annotators were fairly compensated above local minimum wage in accordance with regional labor laws. They received clear instructions on task objectives, scoring criteria, and example cases prior to annotation. No personal or sensitive information was collected, and no individual-level data beyond task performance was recorded.
LLMs like GPT-5 were used appropriately in this work, primarily for language polishing. We ensured that all content was human-authored and carefully reviewed, with no reliance on LLMs for generating scientific claims or experimental results, thereby avoiding potential AI hallucinations.
Appendix H Details of the AuditEval-llm Prompt
Figure 6 and Figure 7 illustrate the prompts used in AuditEval-llm, which instruct the model to act as a professional audio editing quality evaluator.
Prompt You are a professional audio editing quality evaluator. You will be given: (1) the original audio, (2) the edited audio, (3) a textual description of the original audio (original_caption), (4) a textual description of the target edited audio (target_caption). The semantic difference between original_caption and target_caption implicitly defines the intended audio edit. Your task is to evaluate how well the edited audio reflects this semantic change, while maintaining high audio quality and preserving unedited content. Evaluate the edited audio from three perspectives: Quality, Relevance, and Faithfulness. Each score should be an integer from 1 to 5 (higher is better). Output only three integer scores, without explanations. 1. Quality (1-5): Evaluate the perceptual quality of the edited audio compared to the original, including intelligibility, fluency, naturalness, and background noise. - 5: As good as or better than the original, with clear, fluent, and natural audio. - 4: Slight degradation in at most one aspect, but overall quality remains acceptable. - 3: Noticeable degradation in multiple aspects or clear degradation in one aspect, but still understandable.
Figure 6: Examples AuditEval-llm Promps(I)
Prompt - 2: Significant artifacts that seriously harm naturalness or intelligibility. - 1: Severely damaged or unintelligible audio. 2. Relevance (1-5): Evaluate how well the edited audio matches the semantic content described in target_caption, relative to original_caption. Focus only on the parts that are different between original_caption and target_caption. - 5: The edited audio accurately and naturally reflects the target_caption, with correct and complete semantic changes. - 4: The main semantic change is correctly reflected, with only minor imperfections. - 3: The general semantic intent is correct, but details are partially misaligned or awkward. - 2: The semantic change is loosely reflected, with clear errors or unnatural execution. - 1: The edited audio does not match the target_caption. 3. Faithfulness (1-5): Evaluate how well the audio content that should remain unchanged (according to the captions) is preserved in the edited audio, in terms of rhythm, order, loudness, prosody, and style. Do not penalize the semantic changes required by target_caption. - 5: Unchanged content is perfectly preserved and indistinguishable from the original. - 4: Mostly preserved, with only minor variation. - 3: Noticeable changes in rhythm or prosody, but overall style remains similar. - 2: Major changes in unedited content with limited resemblance. - 1: Unedited content is heavily altered or missing. Original Caption: {original_caption} Target Caption: {target_caption}
Figure 7: Examples AuditEval-llm Promps(II)
@@ Appendix H Details of the AuditEval-llm Prompt
Appendix I Evaluation scores Across Editing Operations (Add / Delete / Replace)
This section presents a breakdown of evaluation score across different audio editing operations, including Add, Delete, and Replace, as shown in following table: